The bacterium E. Coli is a Gram-negative rod that is 1mM  x 2mM.   Each E. Coli cell contains...

  1. The bacterium E. Coli is a Gram-negative rod that is 1mM  x 2mM.   Each E. Coli cell contains a chromosome of 4 x 106bp of double-stranded DNA.  Since 1 kilobase pair (kbp) of double-stranded DNA is 0.340mM in length, what is the length of the E. Coli chromosome? (1 kbp = 1000 base pairs, and 1mM = 10-6meters).   
  2. A single colony-forming unit (CFU) of E. coli contains 109 cells.  If you were to join all the E. coli chromosomes contained in that colony together into one gigantic chromosome, how long would it be? Express your answer in miles (1 meter = 6.2 x 10-4mile)

need help with part B!
Answer #1

length of chromosome in basepairs=4 10^6 basepairs.

length of 1000 bp=0.34 mm

so length of 1 bp= 0.34/1000 mm

length of 4 10^6 basepairs.=4 10^6 (0.34/1000)= 1360 mm

length of chromosomes in 109 cells= length of chromosome in one cell 109=1360 10^9=1.3610^12 mm

1 mm=10^-3 m

so 1.3610^12 mm=1.36 10^1210^-3=1.3610^9 m.

1 meter = 6.2 x 10-4mile

so

1.3610^9 m=1.3610^9 6.210^-4= 843200 miles

so the answer is 843200 miles